40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 13 13:35:38 2009, Yoichi, Update, LSC, Locking status 
    Reply  Sat Feb 14 16:53:26 2009, rob, Update, LSC, Locking status 
       Reply  Sun Feb 15 09:35:00 2009, Yoichi, Update, LSC, Locking status 
       Reply  Sun Feb 15 15:53:21 2009, Rob, Update, LSC, Locking status 
          Reply  Mon Feb 16 10:18:13 2009, Alberto, Update, LSC, FE system rebooted 
             Reply  Mon Feb 16 14:12:21 2009, Yoichi, Update, LSC, FE system rebooted 
Message ID: 1301     Entry time: Fri Feb 13 13:35:38 2009     Reply to this: 1304
Author: Yoichi 
Type: Update 
Category: LSC 
Subject: Locking status 
Yoichi, Jenne, Alberto, Rob

Last night, the locking proceeded until the CARM -> MC_L hand-off.
However, the MC_F gets saturated (as expected) and the IFO loses lock soon after the hand-off.
So we need to offload MC_F.
We ran the offloadMCF script, but it did not work, i.e. just waiting for CARM mode.
Looks like an EPICS flag is not set right.
ELOG V3.1.3-